A speedometer reads a speed at 103 km/hr. However, the actual speed the car is going is 98km/hr. What would be the percentage error?
1
Expert's answer
2019-09-17T12:08:08-0400

The error of a speedometer


"\varepsilon=\frac{\Delta v}{v}\times 100\%"

"=\frac{103-98}{98}\times 100\%=5.1\%"